OpenStudy (anonymous):

what are the amplitude phase pellet and period of f(x)=-7sin(4x-pi)+2

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle f(t)=A\sin(\omega t+\phi) }\) \(\color{#000000}{ \displaystyle A }\) - amplitude \(\color{#000000}{ \displaystyle 2\pi/\omega }\) - period \(\color{#000000}{ \displaystyle \phi }\) - phase shift

OpenStudy (solomonzelman):

And in general, whenever you have a parent function, \(\color{#000000}{ \displaystyle y=f(t) }\), and you then add a constant \(c\) to that. i.e., \(\color{#000000}{ \displaystyle y=f(t)+c }\), then the new function is shifted by \(c\) units vertically (if \(c>0\) then up, if \(c<0\) then down).

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

im not good at this at all can you show how its done

OpenStudy (solomonzelman):

I can give you an example if you like. Ok?

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

Actually before posting an example, I want to correct myself. The amplitude is |A| (the absolute value of A, so if A is -5, then amplitude is just 5)

OpenStudy (anonymous):

whats the formula for it

OpenStudy (solomonzelman):

I gave you the formula for it

OpenStudy (anonymous):

a=7?

OpenStudy (anonymous):

oh sorry lol

OpenStudy (solomonzelman):

yes, very good! That would be the amplitude in your case

OpenStudy (solomonzelman):

Can you try the period and the phase shift as well? (I will correct you if anything)

OpenStudy (anonymous):

ok hold on

OpenStudy (solomonzelman):

Sure, take your time:)

OpenStudy (anonymous):

is 2pi over 4x the period

OpenStudy (anonymous):

4x?

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle f(x)=A\sin(\omega x+\phi) }\) has a period \(2\pi/\omega\).

OpenStudy (anonymous):

could you explain to me sorry

OpenStudy (solomonzelman):

so, \(\color{#000000}{ \displaystyle f(x)=-7\sin(4 x-\pi)+2 }\) has a period of \(2\pi/4\)

OpenStudy (solomonzelman):

see how these two are similar?

OpenStudy (anonymous):

yes so thats the period

OpenStudy (solomonzelman):

Yes, and you can simplify that in this case. \(2\pi/4=\pi/2\), since \(2\) and \(4\) can be divisible by 2.

OpenStudy (solomonzelman):

Ok, good with amplitude and period so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i need the phase shift

OpenStudy (solomonzelman):

Ok.

OpenStudy (solomonzelman):

I made a too quick judgement, I have another error in my formula. THis error relates to my form of phase shift.

OpenStudy (anonymous):

oh ok lol

OpenStudy (solomonzelman):

In general, if you have a parent function. \(\color{#000000}{ \displaystyle y=f(x) }\) then to shift it c units right (or \(c\) units left if \(c\) is negative), \(\color{#000000}{ \displaystyle y=f(x-c) }\)

OpenStudy (anonymous):

c is -pi

OpenStudy (solomonzelman):

That means that if you have a function in a form \(\color{#000000}{ \displaystyle y=A\sin(bx) }\) has a phase shift of \(c\) units right, of the following form: \(\color{#000000}{ \displaystyle y=A\sin(b(x-c)) }\) (Not just \(\color{#000000}{ \displaystyle y=A\sin(bx-c) }\) as I previously claimed)

OpenStudy (solomonzelman):

So, in your case, you have: \(\color{#000000}{ \displaystyle y=-7\sin(\color{red}{4x-\pi} ) +2 }\) you will have to re-write the part in red, in a \(b(x-c)\) form (because as it is now, it is in a form of \(bx-c\).) (and that's is very different, right?)

OpenStudy (anonymous):

so pi +4x

OpenStudy (anonymous):

no4(x-pi)

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle 4x-\pi=4\cdot x -4\cdot \frac{\pi}{4}=4\left(x-\frac{\pi}{4}\right) }\)

OpenStudy (anonymous):

so whats the phase shift then? thats confusing lol

OpenStudy (solomonzelman):

So, by factoring I can obtain that \(\color{#000000}{ \displaystyle f(x)=-7\sin(4x-\pi ) +2 }\) is the same thing as \(\color{#000000}{ \displaystyle f(x)=-7\sin\left[4\left(x-\frac{\pi}{4}\right)\right] +2 }\)

OpenStudy (anonymous):

4(x-pi/4)?

OpenStudy (solomonzelman):

remember that y=f(x-c) is phase shifted by c units from y=f(x).

OpenStudy (solomonzelman):

So, this "c" in your case is ?

OpenStudy (anonymous):

pi

OpenStudy (solomonzelman):

try again. \(\color{#000000}{ \displaystyle f(x)=-7\sin\left[4\left(x-\frac{\pi}{4}\right)\right] +2 }\)

OpenStudy (solomonzelman):

What do we subtract from \(x\) ?

OpenStudy (anonymous):

-pi/4

OpenStudy (anonymous):

so thats the shift?

OpenStudy (solomonzelman):

Yes, be we subtract just \(\pi/4\) .

OpenStudy (solomonzelman):

So the phase shift is \(\pi/4\)

OpenStudy (anonymous):

ahhh thank you so much could you help on another one ?

OpenStudy (anonymous):

isnt the pi negat

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

if sin theta =3/4 what is the vale of cos theta?

OpenStudy (solomonzelman):

It might be confusing often, but this is what you need to know. A function in a form \(\color{#000000}{ \displaystyle f(x)=A\sin(\omega[t-\phi]) }\) has an amplitude |A|, phase shift \(\phi\), and a period \(2\pi/\omega\)

OpenStudy (anonymous):

ohhh ok

OpenStudy (solomonzelman):

No it is not negative.

OpenStudy (anonymous):

thank you could you help on the one above please

OpenStudy (solomonzelman):

\(\sin \theta = \frac{3}{4}\)

OpenStudy (solomonzelman):

sin = opp / hyp correct?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

Yes, so using the pythagorean theorem (\(a^2+b^2=c^2\)), can you find the adjacent side? (Hint: you know that \(\rm Opposite~Side=3\), and \(\rm Hypotenuse=4\))

OpenStudy (anonymous):

so c^2=3^2+4^2

OpenStudy (solomonzelman):

Be careful, although nice try. "c" is always the hypotenuse! So, you should rather say, \(3^2+b^2=4^2\)

OpenStudy (anonymous):

oh ok

OpenStudy (solomonzelman):

Ok, can you find \(b\) ?

OpenStudy (anonymous):

9+b^2=16 subtract 9

OpenStudy (anonymous):

b^=7

OpenStudy (solomonzelman):

you mean \(b^2=7\). Right?

OpenStudy (anonymous):

yea

OpenStudy (solomonzelman):

And if \(b^2=7\), then \(b=?\)

OpenStudy (solomonzelman):

(Hint: Take the square root of both sides)

OpenStudy (anonymous):

b=

OpenStudy (solomonzelman):

If you want to type a square root of a number c, you can use: `\(\sqrt{~ c ~} \)` (~ is ajust a space in latex, and you can copy the code).

OpenStudy (anonymous):

2.64

OpenStudy (solomonzelman):

2.6\(\color{red}{45}\)7 --> rounds to 2.6\(\color{blue}{5}\)

OpenStudy (anonymous):

ok thank you

OpenStudy (solomonzelman):

So, if you want to approximate the square root, instead of leaving it as an exact number, you can certainly put \(2.65\).

OpenStudy (solomonzelman):

We have only found the adjacent side.

OpenStudy (anonymous):

oh

OpenStudy (solomonzelman):

Now you know that the adjacent side is 2.65 and you were given that the hypotenuse is 4. So, knowing that \(\cos \theta = \displaystyle \frac{\rm adjacent~side}{\rm ~hypotenuse} \)

OpenStudy (anonymous):

2.65/4?

OpenStudy (solomonzelman):

you will therefore get that \(\cos \theta = ?~\) yes, and that is correct

OpenStudy (anonymous):

0.6625

OpenStudy (solomonzelman):

you can re-write it otherwise.... yes, that would also suffice :)

OpenStudy (anonymous):

thank you so much

OpenStudy (solomonzelman):

Not a problem:)

OpenStudy (mathmale):

@areeb50: Next time, please leave the word "pellet" out of your problem statement. What did it gain you to include that the first time?

OpenStudy (anonymous):

@mathmale it was an accident ._. english isnt my first language